OpenStudy (allymariex3):

Can anyone explain Factoring by Grouping (four-term polynomials and trinomials) please and thank you!

OpenStudy (anonymous):

Let's use a concrete example. Consider the following trinomial: 2x^2 + 7x + 6 To factor by grouping, we need at least four terms. We can split up the middle term to accomplish this: 2x^2 + 4x + 3x + 6 = 2x(x+2) + 3(x + 2) = (x+2)(2x+3) Hopefully this will give you a little insight into how to change trinomials into four-term polynomials that can be factored by grouping. Even I'm not certain how to explain the question fully, but I hope this can get you started.